A 17.0-g sample of HF is dissolved in water to give 200. mL of solution. The concentration of the solution is:
Select one:
a. 0.17 M
b. 8.5 M
c. 4.2 M
d. 0.85 M
e. 0.09 M
Answer #1

Molar mass of HF=20.01g/mol

Mole=17/20.01=0.85 mole hf

Concentration= mole/volume in litre

Volume=0.2 litre

Concentration=0.85/0.2= 4.2 M

Answer C